An automobile tire is inflated to 32 psig pressure at 50 F. After being driven the temperature rise to 75F. Determine the final gage pressure assuming the volume remains constant.


1
Expert's answer
2021-11-29T11:38:41-0500

Process is isochoric

Ideal Gas Law

"pV=nRT \\

pV=RT \\

\frac{p}{T} = \frac{R}{V} = constant \\

\frac{p_1}{T_1} = \frac{p_2}{T_2} \\

p_2 = \frac{p_1 \times T_2}{T_1} \\

T_2 = 75 \; F = 297 \;K \\

T_1 = 50 \;F = 283 \;K \\

p_2 = \frac{32 \times 297}{283} = 33.58 \;psig"